OpenStudy (anonymous):

How to divide problems like the following: (y+2)/(y+6) ....... don't know where to start, I know the answer but I don't know how to get it myself. Even a helpful video link will suffice.

OpenStudy (anonymous):

the numerator and denominator both have degree 1, so you will get a number, plus a remainder there is an easy trick for this would you like to see it?

OpenStudy (anonymous):

yes please !

OpenStudy (anonymous):

in other words, you know when you divide it will be \[\frac{y+2}{y+6}=1+\frac{\text{some number}}{y+6}\]

OpenStudy (anonymous):

the only question is, what is the number up top since you know \(y\) goes in to \(y\) one time here is the trick

OpenStudy (anonymous):

\[\frac{y+2}{y+6}=\frac{y+6-4}{y+6}=\frac{y+6}{y+6}-\frac{4}{y+6}=1-\frac{4}{y+6}\]

OpenStudy (anonymous):

you can do it mentally pretty much, i just wrote out the steps to show what i was doing

OpenStudy (anonymous):

you can also do long division but a) it takes longer and b) i can't write long division here

OpenStudy (anonymous):

Oh i see, but why 4 ?

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

the denominator is \(y+6\) so i forced the numerator to be \(y+6\) as well, by adding and subtracting \(4\) up top

OpenStudy (anonymous):

i wrote \[y+2\] as \[y+6-4\] in other wrods

OpenStudy (anonymous):

that way i get the 1 i need \[\frac{y+6}{y+6}=1\] and the remainder is \(-4\)

OpenStudy (anonymous):

I think i get what you mean.

OpenStudy (anonymous):

Thank you!